# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v4

@@ Abstract
We introduce the Cooperative Network Architecture (CNA), a model that represents sensory signals using structured, recurrently connected networks of neurons, termed “nets.” Nets are dynamically assembled from overlapping net fragments, which are learned based on statistical regularities in sensory input. This architecture offers robustness to noise, deformation, and out-of-distribution data, addressing challenges in current vision systems from a novel perspective. We demonstrate that net fragments can be learned without supervision and flexibly recombined to encode novel patterns, enabling figure completion and resilience to noise. Our findings establish CNA as a promising paradigm for developing neural representations that integrate local feature processing with global structure formation, providing a foundation for future research on invariant object recognition.
@@ 1 Introduction
Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Sager et al., 2025). Yet, despite these advances, artificial vision systems continue to struggle with robust object representation and recognition, particularly under varying viewing conditions, occlusion, or noise (Qi et al., 2023; Simmler et al., 2021). This limitation highlights a fundamental gap between current AI systems and the flexible, context-aware perception observed in biological vision.
@@ 2 The Net Fragment Framework
Additionally, we limit the weights to a range between 0 and 111, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
@@
The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment. We call this accumulated input the pre-synaptic activity at timestep $t$, which is a function $\mathcal{I}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
@@
Neurons are not only influenced by direct connections within net fragments but also by indirect connections (e.g., if $n_{i}$ supports $n_{j}$ and $n_{j}$ supports $n_{k}$, then $n_{i}$ indirectly supports $n_{k}$). A set of directly and indirectly connected neurons is called a net, implementing the basic structure used in this work to represent objects. For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active as well:
@@ 3.1 Stage 1: Feature Extraction
The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels (111 for grayscale, 333 for RGB), while $H,W$ stand for the height and width of the image.
@@ 3.2 Stage 2: Net Fragments
$\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)$ =$=\left\{n_{j}^{(S1)}\mid d(n_{i}^{(S2)},n_{j}^{(S1)})\leq r\right\}$
@@
The pre-synaptic activity is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{I}},\mathcal{E}\right)$. In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{I}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range [0,1][0,1][ 0 , 1 ] (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active. Specifically, attenuation is implemented as:
@@
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all competitive $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to 111 if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of competitive neurons. The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to 111.
@@ 4.3 Compositionality
Figure 4: Training and evaluation data. The top of the figure illustrates the straight lines used during the training phase. Below, various samples from the evaluation datasets are shown, including kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888). For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
The model weights are trained using straight lines (see training data examples at the top of Figure 4). During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 111888 of Figure 4). Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows 111 and 222), digits and characters (rows 333 and 444), and line drawings (rows 555888).
@@ 4.4 Filtering Gaussian Noise
[…]
@@ 4.5 Reconstructing Figures
Figure 7: Recovery from occlusion. Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$. Part B: A diagonal line with 666 missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the pre-synaptic activity $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
If a net fragment $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$). Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns. This reconstruction is reliable for up to 333 missing pixels (see Appendix G). In some cases, as in this figure, the line is fully reconstructed for up to 666 removed pixels and partly reconstructed for 777 pixels.
@@ 4.6 Comparison to Autoencoders
Figure 8: Comparison of the outputs generated by the proposed CNA model and an autoencoder. Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder). Rows 111333 illustrate how both models encode patterns that were not present in the training dataset. Rows 444555 let one compare the models’ performance in denoising tasks, while row 666 displays their ability to reconstruct lines.
Figure 8 compares performance between the proposed CNA model and an autoencoder (for model and training details, refer to Appendix F). While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training. For instance, rows 111333 demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
[…]
In row 666, the models’ capabilities for reconstructing missing line segments are compared. Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed. This observation is confirmed by the metrics in Appendix G: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of increased noise elsewhere, leading to lower overall precision and recall for figure reconstruction.
@@ 5.2 Boltzmann Machine
The BM as described in Hinton et al. (1984) and our model differ, however, in a number of ways. Inspired by statistical mechanics, the BM speaks of equilibrium macro-states, each of which is realized as an ensemble of micro-states, the probability of a micro-state being proportional to a negative exponential of a state “energy” (Boltzmann distribution), which is defined as in Amari-Hopfield’s Associative Memory as a function of neural activities and connection weights. The BM further assumes that the system compares two modes, clamped (in which the sensory neurons are determined by the input) and free running (no sensory input), and lets synaptic plasticity equalize the signal correlations of pairs of neurons observed in one of the two modes with those observed in the other. In our model, neural activity is not modeled in terms of probabilistic ensembles of states but as discrete-time trajectories over a fixed number of time steps $T$, during which the same sensory input is constantly refined. These trajectories evolve deterministically according to update rules that depend on both the input and recurrent connectivity. Importantly, the dynamics are externally modulated: At each step, a control mechanism attenuates the activity of neurons that lack sufficient recurrent support from other active neurons; see Eqs. 28 and 29. In contrast, Boltzmann Machines rely on stochastic sampling over ensembles of network states, governed by a distribution derived from an energy function. While the statistical mechanics framework underlying Boltzmann Machines is elegant and conceptually rich, it is computationally inefficient in settings where a large number of intermediate states must be generated and compared to approximate equilibrium. Our model avoids this bottleneck by converging rapidly through deterministic dynamics over a small fixed number of steps.
@@
The realization that the BM in its generality would imply unfeasible learning times led to the introduction of Restricted Boltzmann Machines (RBM) (Smolenksy, 1987; Hinton and Salakhutdinov, 2006). RBMs consist of two layers, one visible and one latent, with bidirectional connections between them, but no connections within them. A stack of RBMs is used in (Hinton and Salakhutdinov, 2006) to construct an autoencoder in which RBM learning is used to pre-train the connections between layers in a sequence from bottom (input) to top (the sparse coding layer). Then, the whole encoder-decoder stack is trained end-to-end by backpropagation of error. We discuss the differences between our model and networks trained by backpropagation of error (including RBMs) in the next sub-section.
@@ Appendix A Feature Extraction
The stage $S1$ serves to extract initial binary features from the images. The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $\times 5$ 5 as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (‘stride’) of 111 (assuming 0 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size (1×32×32)(1$\times$ 32$\times$ 32)( 1 × 32 × 32 ) and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size (4×32×32)(4$\times$ 32$\times$ 32)( 4 × 32 × 32 ). The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using the heaviside function $\Theta$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 18).
@@ Appendix B Normalization
The pre-synaptic activity in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{I}(S2)}$ to the interval (0,1)(0,1)( 0 , 1 ).
@@
According to this formula, $\boldsymbol{a}^{\mathcal{I}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{I}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope 1/2-1/2- 1 / 2. (In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well). This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons. In the next step, we introduce the normalization
@@
Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$. This normalization confines activations to the interval [0,1][0,1][ 0 , 1 ], where neurons receiving maximal support are mapped to 111, while those with less support are mapped to correspondingly smaller values. This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.

### v5

@@ Abstract
We introduce the Cooperative Network Architecture (CNA), a model that represents sensory signals using structured, recurrently connected networks of neurons, termed “nets.” Nets are dynamically assembled from overlapping net fragments, which are learned based on statistical regularities in sensory input. This architecture offers robustness to noise, deformation, and generalization to out-of-distribution data, addressing challenges in current vision systems from a novel perspective. We demonstrate that net fragments can be learned without supervision and flexibly recombined to encode novel patterns, enabling figure completion and resilience to noise. Our findings establish CNA as a promising paradigm for developing neural representations that integrate local feature processing with global structure formation, providing a foundation for future research on invariant object recognition.
@@ 1 Introduction
Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Sager et al., 2025). Yet, despite these advances, artificial vision systems continue to struggle with robust object representation and recognition, particularly under varying viewing conditions, occlusion, or noise (Fan et al., 2023; Simmler et al., 2021). This limitation highlights a fundamental gap between current AI systems and the flexible, context-aware perception observed in biological vision.
@@ 2 The Net Fragment Framework
Additionally, we limit the weights to a range between 0 and $1$, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
@@
The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment. We call this accumulated input the pre-synaptic activity at timestep $t$, which is a function $\mathcal{I}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep t1t-1.
@@
Neurons are not only influenced by direct connections within net fragments but also by indirect connections (e.g., if $n_{i}$ supports $n_{j}$ and $n_{j}$ supports $n_{k}$, then $n_{i}$ indirectly supports $n_{k}$). A set of directly and indirectly supported neurons is called a net, implementing the basic structure used in this work to represent objects. For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active as well:
@@ 3.1 Stage 1: Feature Extraction
The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels ($1$ for grayscale, $3$ for RGB), while $H,W$ stand for the height and width of the image.
@@ 3.2 Stage 2: Net Fragments
$\mathcal{B}_{r}\left(n^{(S1)}_{i}\right)$ $=\left\{n_{j}^{(S1)}\mid d(n_{i}^{(S2)},n_{j}^{(S1)})\leq r\right\}$
@@
The pre-synaptic activity is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{I}},\mathcal{E}\right)$. In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{I}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range $[0,1]$ (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active. Specifically, attenuation is implemented as:
@@
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all competitive $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to $1$ if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of competitive neurons. The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to $1$.
@@ 4.3 Compositionality
Figure 4: Training and evaluation data. The top of the figure illustrates the straight lines used during the training phase. Below, various samples from the evaluation datasets are shown, including kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows 5588). For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
The model weights are trained using straight lines (see training data examples at the top of Figure 4). During inference – without additional learning – the learned net fragments are recursively composed to form nets representing significantly more complex structures (see validation data in rows 1188 of Figure 4). Notably, although such structures are absent from the training set, the CNA successfully generalizes to represent kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows 5588).
@@ 4.4 Filtering Gaussian Noise
To further investigate noise filtering, we introduce Gaussian noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - resulting from adding up to $20\%$ noise per channel across four independent channels; see Appendix D) to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
The results are visualized in Figure 5: While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated. As additional noise is introduced, more undesired neurons receive support and persist in their activity. However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$. In Appendix G, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
Figure 6: The filtering of (simulated) noise over time: Already at time step $t=0$, most of the noise is suppressed. However, when attenuation increases over time, even more noise is filtered out. The plot is created after adding $10\%$ noise to each feature channel (using a bias $b^{(S2)}=0.7$).
@@ 4.5 Reconstructing Figures
Figure 7: Recovery from occlusion. Part A: Input images of diagonal lines with different numbers of missing pixels, corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, attenuation coefficient $\gamma=0.6+0.2t$. Part B: A diagonal line with 66 missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the pre-synaptic activity $\boldsymbol{a}^{(S2)}$ displaying the “activation probabilities” (on the right) after the first time step.
If a net fragment $\mathcal{F}_{t}(n_{i})$ contains a substantial number of (active) neurons, an inactive neuron $n_{i}$ can receive significant support, leading it to switch on and encouraging figure reconstruction ($\mathcal{F}_{t}(n_{i})\rightarrow\mathcal{F}^{*}_{t}(n_{i})$). Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns. This reconstruction is reliable for up to $3$ missing pixels (see Appendix G). In some cases, as in this figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.
@@ 4.6 Comparison to Autoencoders
Figure 8: Comparison of the outputs generated by the proposed CNA model and an autoencoder. Each row presents two samples, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, alongside their respective feature activations and resulting representations (once for the CNA model and once for the autoencoder). Rows 1133 illustrate how both models encode patterns that were not present in the training dataset. Rows 4455 let one compare the models’ performance in denoising tasks, while row 66 displays their ability to reconstruct lines.
Figure 8 compares performance between the proposed CNA model and an autoencoder (for model and training details, refer to Appendix F). While the autoencoder successfully learns to accurately represent line structures observed in the training data – achieving a mean square error (MSE) of less than $1\times 10^{-3}$ between the input and the reconstruction by the decoder – it struggles with generalization to data not observed during training. For instance, rows 1133 demonstrate that the CNA model is capable of representing previously unseen structures, while, in contrast, the autoencoder fails to capture such out-of-distribution patterns.
Rows 4455 highlight the autoencoder’s ability to represent line structures with up to 11.5% noise introduced. However, as the noise level increases, its performance decreases significantly. The CNA model, in contrast, remains robust under much higher noise levels, tolerating up to $59\%$ noise while still producing high-quality representations. Metrics on noise reduction, precision, and recall, presented in Appendix G, further substantiate the CNA model’s superior performance.
In row $6$, the models’ capabilities for reconstructing missing line segments are compared. Both models can reconstruct the removed pixels, but the autoencoder introduces additional noise in regions other than the line center, where the pixels were removed. This observation is confirmed by the metrics in Appendix G: The autoencoder is slightly superior in reconstructing the removed pixels, but this comes at the cost of increased noise elsewhere, leading to lower overall precision and recall for figure reconstruction.
@@ 5.2 Boltzmann Machine
The BM as described in Hinton et al. (1984) and our model differ, however, in a number of ways. Inspired by statistical mechanics, the BM speaks of equilibrium macro-states, each of which is realized as an ensemble of micro-states, the probability of a micro-state being proportional to a negative exponential of a state “energy” (Boltzmann distribution), which is defined as in Amari-Hopfield’s Associative Memory as a function of neural activities and connection weights. The BM further assumes that the system compares two modes, clamped (in which the sensory neurons are determined by the input) and free running (no sensory input), and lets synaptic plasticity equalize the signal correlations of pairs of neurons observed in one of the two modes with those observed in the other. In our model, neural activity is not modeled in terms of probabilistic ensembles of states but as discrete-time trajectories over a fixed number of time steps $T$, during which the same sensory input is constantly refined. In distinction to the Boltzmann Machine, these trajectories evolve deterministically, according to update rules that depend on both the input and recurrent connectivity. Importantly, the dynamics are externally modulated: At each step, a control mechanism attenuates the activity of neurons that lack sufficient recurrent support from other active neurons; see Eqs. 28 and 29. In contrast, Boltzmann Machines rely on stochastic sampling over ensembles of network states, governed by a distribution derived from an energy function. While the statistical mechanics framework underlying Boltzmann Machines is elegant and conceptually rich, it is computationally inefficient in settings where a large number of intermediate states must be generated and compared to approximate equilibrium. Our model avoids this bottleneck by converging rapidly through deterministic dynamics over a small fixed number of steps.
@@
The realization that the BM in its generality would imply unfeasible learning times led to the introduction of Restricted Boltzmann Machines (RBM) (Smolensky, 1987; Hinton and Salakhutdinov, 2006). RBMs consist of two layers, one visible and one latent, with bidirectional connections between them, but no connections within them. A stack of RBMs is used in (Hinton and Salakhutdinov, 2006) to construct an autoencoder in which RBM learning is used to pre-train the connections between layers in a sequence from bottom (input) to top (the sparse coding layer). Then, the whole encoder-decoder stack is trained end-to-end by backpropagation of error. We discuss the differences between our model and networks trained by backpropagation of error (including RBMs) in the next sub-section.
@@ Appendix A Feature Extraction
The stage $S1$ serves to extract initial binary features from the images. The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that straight lines when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (‘stride’) of $1$ (assuming 0 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$. The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using the heaviside function $\Theta$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 18).
@@ Appendix B Normalization
The pre-synaptic activity in $S2$ is normalized using a two-step approach that confines the neuronal activations $\boldsymbol{a}^{\mathcal{I}(S2)}$ to the interval $(0,1)$.
@@
According to this formula, $\boldsymbol{a}^{\mathcal{I}}$ grows undiminished with $\boldsymbol{a}^{\mathcal{I}}$ until it reaches the value $\lambda$, and thereafter actually diminishes with slope 1/2-1/2. (In our experiments we found the value $\lambda=1.3\cdot\frac{h+w}{2}$ to work well). This saturation helps to mitigate imbalances between net fragments with differing numbers of participating neurons. In the next step, we introduce the normalization
@@
Here, $i$ refers to spatial locations within the image, i.e., $i=(h,w)$ where $h\in H,w\in W$. This normalization confines activations to the interval $[0,1]$, where neurons receiving maximal support are mapped to $1$, while those with less support are mapped to correspondingly smaller values. This normalization is relative to the maximal value that itself is growing during learning as it is important to give neurons initially (when recurrent connections are still small) a chance to fire and later when recurrent weights between correlated patterns have been learned, to suppress those neurons that are part of noise patterns.